OpenStudy (anonymous):

A rectangular prism has length , 13 m width 6 m , and volume 390m3 . What is the height of the prism?

OpenStudy (anonymous):

Are you going to actually do your own homework?

OpenStudy (anonymous):

V= w l h

OpenStudy (anonymous):

\[ h = \frac V{w l}=\frac {390}{6\times 13}=5 \]